# Robust Containment Queries over Collections of Rational Parametric Curves via Generalized Winding Numbers

### v1

@@ 7.2. Algorithm Performance
One explanation for the relatively poor performance of the two ray casting algorithms is that such procedures are unnecessarily informative. Not only do ray casting algorithms identify potential intersections, but they also inherently locate the intersections, information which is rarely needed when the ray itself is chosen arbitrarily. In contrast, by not needing to compute the location of these intersections, the proposed algorithm is able to converge with considerably fewer curve evaluations, despite only doing so linearly. This is particularly relevant in the case when the curves of interest are rational and contain significant curvature, multiple overlapping components, and self-intersections, as in Figure 17. In these examples, the ray casting algorithms perform particularly poorly, as any ray extended will intersect the curve multiple times, incurring additional cost with each.
@@
Figure 17. For higher-order curves, additional curve subdivisions are necessary to reach perfect geometric fidelity. The proposed algorithm still has superior performance in these examples, as the location of the various intersections need not be computed directly.
@@
As is always the case when generalized winding numbers are used to determine containment, there is still an assumption that the collection of curves are properly oriented. Because the scalar field generated by a single curve is computed completely independently of every other curve by design, a single reversed curve interferes with the contribution of surrounding curves, functionally reversing nearby containment queries, as can be observed in Figure 18. While the induced classification errors are still local to the geometric error, and would no doubt cause catastrophic errors without the use of generalized winding numbers, they nevertheless represent a type of geometric error unaccounted for by the current method.
@@ 8. Concluding remarks
In this paper, we proposed a generalized winding number algorithm for collections of parametric curves, addressing the inherent challenges for curved geometry that are not shared by their linear counterparts. For the vast majority of points that are considered “sufficiently far” from a given curve, we circumvent the need for expensive linearization of the shape by treating the curve directly as a single segment connecting its endpoints. For all other points, we demonstrated that our recursive algorithm outperforms existing ray casting algorithms adapted for this context. Finally, we formalize a procedure for handling winding numbers of points that are coincident with the curve, further increasing the robustness of our algorithm. In each case, this algorithm makes minimal reference to the often near-singular integrals from which the winding number is theoretically derived, instead operating strictly according to geometric and trigonometric principles. Nevertheless, we are able to exactly compute generalized numbers for curved shapes in two dimensions. Whereas the focus of this work has been on efficiently and robustly computing winding numbers for individual curves within a collection, follow up work will consider accelerating the overall query workflow using a spatial index, as in Jacobson et al. (2013); Barill et al. (2018); Weiss et al. (2016) and exploiting the inherent parallelism within Algorithm 1 via threaded and/or GPU implementation.
@@
Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344.

### v2

@@ 7.2. Algorithm Performance
One explanation for the relatively poor performance of the two ray casting algorithms is that such procedures are unnecessarily informative. Not only do ray casting algorithms identify potential intersections, but they also inherently locate the intersections, information which is rarely needed when the ray itself is chosen arbitrarily. In contrast, by not needing to compute the location of these intersections, the proposed algorithm is able to converge with considerably fewer curve evaluations, despite only doing so linearly. This is particularly relevant in the case where the curves of interest contain multiple overlapping components, self-intersections, or regions of significant curvature, the latter of these cases being tested directly on a rational curve in Figure 16. In these examples, the ray casting algorithms perform particularly poorly, as any ray extended will intersect the curve multiple times, incurring additional cost with each.
@@
Figure 16. For higher-order curves, additional curve subdivisions are necessary to reach perfect geometric fidelity. The proposed algorithm still has superior performance in such cases, as the location of the various intersections need not be computed directly.
[…]
Figure 16: An identical plot to Figure 15 is shown, but on a high-order rational Bezier curve in the shape of a spiral. While all three methods need more evaluations than for the simpler watertight shape, the proposed method still requires the fewest curve evaluations for the vast majority of points.
Finally, we consider the sensitivity of our procedure for computing generalized winding numbers to the numerical tolerance introduced in Algorithm 3. We do this by considering performance and accuracy as a function of this tolerance, respectively measured in terms of the number of curve evaluations required to determine containment and the number of misclassifications caused by the final layer of implied linearization. These metrics are evaluated over $10^{5}$ points sampled from a bounding box for the closed shape in Figure 17, which for demonstration purposes depicts a $9^{\text{th}}$ order, rational curve that far exceeds the complexity of those typically used in practice. We see that invoking Algorithm 3 to prematurely terminate the procedure does marginally reduce the overall computational cost, but that this performance improvement is closely correlated with losses in classification accuracy. This effect is amplified on shapes with such complex curvature, as it becomes more likely that the implied linearization imposed by the tolerance will produce misclassifications. At the same time, we observe that even for very intricate shapes, the performance of Algorithm 2 rapidly stabilizes along both metrics. As a result, our practical preference is to set this tolerance to zero even in cases where geometric accuracy is not a priority.
Figure 17. Beyond a certain threshold, the adaptive nature of the proposed algorithm causes it to be insensitive to the numerical tolerance in Algorithm 3. In practice, we set this tolerance to zero, as this ensures exact geometric fidelity while not incurring a large additional computational burden.
[…]
Figure 17: Two bar plots show the number of curve evaluations and misclassifications for the proposed method as a function of the numerical tolerance for the Algorithm 3. These results are reported for randomly sampled points around a high-order rational Bezier curve. When the tolerance is set lower than 10 to the power of -6, the proposed method uses a fixed number of curve evaluations and does not introduce any misclassifications.
@@
As is always the case when generalized winding numbers are used to determine containment, there is still an assumption that the collection of curves are reasonably oriented. Because the scalar field generated by a single curve is computed completely independently of every other curve by design, a single reversed curve interferes with the contribution of surrounding curves, functionally reversing nearby containment queries, as can be observed in Figure 18. While the induced classification errors are still local to the geometric error, and would no doubt cause catastrophic errors without the use of generalized winding numbers, they nevertheless represent a type of geometric error unaccounted for by the current method.
@@
[…]
Figure 18: The same butterfly graphic from Figure 11 is shown, but one curve has the orientation reversed instead of deleted. The reversal causes the winding number field to be reversed in a circular region around the curve.
@@ 8. Concluding remarks
In this paper, we proposed a generalized winding number algorithm for collections of parametric curves, addressing the inherent challenges for curved geometry that are not shared by their linear counterparts. For the vast majority of points that are considered “sufficiently far” from a given curve, we circumvent the need for expensive linearization of the shape by treating the curve directly as a single segment connecting its endpoints. For all other points, we demonstrated that our recursive algorithm outperforms existing ray casting algorithms adapted for this context. Finally, we formalize a procedure for handling winding numbers of points that are coincident with the curve, further increasing the practicality of our algorithm. In each case, this algorithm makes minimal reference to the often near-singular integrals from which the winding number is theoretically derived, instead operating strictly according to geometric and trigonometric principles. Nevertheless, we are able to exactly compute generalized numbers for curved shapes in two dimensions. Whereas the focus of this work has been on efficiently and robustly computing winding numbers for individual curves within a collection, follow-up work will consider accelerating the overall query workflow using a spatial index, as in Jacobson et al. (2013); Barill et al. (2018); Weiss et al. (2016) and exploiting the inherent parallelism within Algorithm 1 via threaded and/or GPU implementation.
@@
Acknowledgements.
We would like to thank the anonymous reviewers for their thoughtful consideration and insightful feedback. This work performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. LLNL-TR-828702.